# ADS-Edit: A Multimodal Knowledge Editing Dataset for Autonomous Driving Systems

### v1

@@ 5.2 Large Multi-modal Models for Autonomous Driving
The recent rise of large multimodal models (LMMs) in autonomous driving (Qian et al., 2024; Cui et al., 2024) has spurred the development of advanced frameworks that integrate perception, prediction, and planning. CODA-LM (Li et al., 2024c) and DriveLM (Sima et al., 2024), have attempted to bridge this gap by integrating hierarchical reasoning and graph-structured visual QA frameworks. CODA-LM evaluates vision-language models on corner cases across perception, prediction, and planning stages, while DriveLM explicitly models logical dependencies between driving phases. Meanwhile, LingoQA (Marcu et al., 2024) expands free-form QA capabilities by integrating action justification and scene understanding. However, these models still face challenges in real-world deployment due to static knowledge representations and modality imbalance. To address these issues, knowledge editing (Yao et al., 2023) has emerged as a promising direction, enabling targeted modifications to model behavior without full retraining. Yet, existing datasets and methods fail to account for the unique demands of autonomous driving, such as multimodal coherence and lifelong adaptation to evolving scenarios. This gap motivates our work, ADS-Edit, the first multimodal knowledge editing dataset designed explicitly for autonomous driving systems.
6 Conclusion
@@
In this paper, we present to leverage knowledge editing techniques to address the challenges faced by Large Multimodal Models in autonomous driving scenarios. ADS-Edit, a new benchmark specifically is designed for evaluating knowledge editing methods in this domain. Through extensive experimental analysis, we systematically compare and evaluate the effectiveness of various editing methods, while providing in-depth insights into the underlying reasons for the observed performance variations and failure cases.
Limitations
While this study advances Knowledge Editing for Autonomous Driving System applications, several limitations warrant consideration:
(1) Only VQA Data.
In this work, we only test knowledge editing baselines on visual question answering data and don’t explore other forms of data related to autonomous driving tasks, such as trajectory prediction. How to successfully update predicted location coordinates as knowledge into LMMs and evaluate their generalization will be explored in future work.
(2) Limited Experimental Baselines.
Due to the high cost of LMMs, we did not conduct experiments on larger scale LMMs and discard knowledge editing methods that require more resources, such as MEND and SERAC.
@@ Appendix B Baselines
WISE (Wang et al., 2024b) uses a dual-memory architecture comprising a main memory for pretrained knowledge and a side memory for editing. A routing mechanism decides which memory to use during inference. Additionally, knowledge is partitioned into multiple subspaces through sharding to facilitate conflict-free edits, with subsequent merging of the shards into a unified side memory.
Appendix C Benchmark Construction Details
Appendix D Case of ADS-Edit benchmark

### v2

@@ 5.2. Large Multi-modal Models for Autonomous Driving
The recent rise of large multimodal models (LMMs) in autonomous driving (Qian et al., 2024; Cui et al., 2024) has spurred the development of advanced frameworks. CODA-LM (Li et al., 2024c) and DriveLM (Sima et al., 2024) integrate hierarchical reasoning and graph-structured visual QA frameworks. Meanwhile, LingoQA (Marcu et al., 2024) expands free-form QA capabilities by integrating action justification and scene understanding. However, these models still face challenges in real-world deployment due to static knowledge representations and modality imbalance.
6. Conclusion
@@
In this paper, we propose leveraging knowledge editing techniques to address the challenges faced by Large Multimodal Models in autonomous driving scenarios. To this end, we introduce ADS-Edit, a novel benchmark specifically designed to evaluate knowledge editing methods within this domain. Through extensive experimental analysis, we systematically compare and assess the effectiveness of various editing approaches, providing comprehensive insights into the underlying causes of performance variations and failure cases. Moreover, our findings highlight that developing editing methods that are both well-balanced and high-performing remains a significant challenge in the autonomous driving systems.
@@ Appendix B Baselines
WISE (Wang et al., 2024d) uses a dual-memory architecture comprising a main memory for pretrained knowledge and a side memory for editing. A routing mechanism decides which memory to use during inference. Additionally, knowledge is partitioned into multiple subspaces through sharding to facilitate conflict-free edits, with subsequent merging of the shards into a unified side memory.
Appendix C Lifelong Editing Results
Appendix D Benchmark Construction Details
Appendix E Case of ADS-Edit benchmark